# Exploiting synergies between JWST and cosmic 21-cm observations to uncover star formation in the early Universe

### v1

@@ 4.3 Joint inference
[…]
@@
Figure 4: Marginal 1D and 2D posteriors of astrophysical parameters from the joint analysis and individual analysis of the datasets: UVLF, CXB, CRB, HERA, and SARAS 3 (color-coded as shown in the legend). The 2D posterior contours show the 68% (dash-dotted) and 95% (dashed) credible regions emerging from the joint analysis. The 1D marginal posterior PDFs show the weighted mean (solid line) and 68% credible region (shaded with dash-dotted outline; calculated as the smallest CDF intervals containing the required probability volume) for the joint fit (black), as well as the 1D PDFs for each individual observable (indicated in the legend). Top right panel shows the posterior to prior volume contraction for each dataset, calculated using margarine. For a similar plot comparing joint constraints with and without the UVLF dataset, see Figure 12. For the full parameter space, including SARAS 3 nuisance parameters, see Figure 13.
Table 3: Marginal 1D constraints on the parameter $V_{c}$, and related quantities $T_{\text{vir}}$ (Equation 5) and $M_{\text{crit}}$ (Equation 3 in the absence of $f_{v_{\text{bc}}}$ and $f_{\text{LW}}$). The quoted errors are at 68% (95%) confidence levels around the weighted mean. We list SARAS 3 and UVLF individually since they are the most constraining datasets, providing lower and upper limits respectively. The joint analysis combining all datasets is, thus, a measurement with this squeezed constraint.
Parameter Prior SARAS 3 UVLF Joint w/ UVLF
$\log_{10}\left(V_{c}/\,\mathrm{km}\,\mathrm{s}^{-1}\right)$ $\mathcal{U}\left(0.62,2\right)$ or $\gtrsim 1.16\,(0.72)$ $\lesssim 1.46\,(1.77)$ $1.23^{+0.40}_{-0.36}\left({}^{+0.56}_{-0.57}\right)$
$\implies V_{c}\ (\,\mathrm{km}\,\mathrm{s}^{-1})$ $\log\mathcal{U}(4.2,100)$ $\gtrsim 14.4\,(5.3)$ $\lesssim 28.7\,(58.2)$ $16.9^{+25.7}_{-9.5}\left({}^{+44.4}_{-12.3}\right)$
$\log_{10}\left(T_{\text{vir}}/\,\text{K}\right)$ $\mathcal{U}\left(3.1,5.9\right)$ $\gtrsim 4.19\,(3.31)$ $\lesssim 4.78\,(5.40)$ $4.32^{+0.80}_{-0.72}\left({}^{+1.12}_{-1.14}\right)$
$\log_{10}\left(M_{\text{crit}}(z=6)/M_{\odot}\right)$ $\mathcal{U}\left(6.5,10.6\right)$ $\gtrsim 8.09\,(6.77)$ $\lesssim 8.98\,(9.90)$ $8.29^{+1.21}_{-1.08}\left({}^{+1.68}_{-1.71}\right)$
@@ 5 Results and discussion
Figure 4 shows the posterior on the astrophysical parameters recovered for the different datasets independently and when they are combined in a joint analysis. A detailed comparison of constraints between the datasets, excluding UVLF observations, is done in Pochinda et al. (2024) using an SFE model that is constant in the atomic cooling regime (dashed orange line in Figure 1). Their results are retrieved in our analysis as a special case of our more general SFE model — i.e., the posterior PDF of $M_{0}$ saturates to the upper prior limit $10^{11}\,\mathrm{M}_{\odot}$, rendering $\alpha_{\star}$ and $\beta_{\star}$ parameters are redundant (since most halos are $<10^{11}\,\mathrm{M}_{\odot}$ at $z\geq 6$). Furthermore, the $\tilde{f}_{\star,\mathrm{II}}$ posterior PDF saturates to the lower prior limit at $10^{-4}$. This is shown in Figure 12, where we compare the joint constraints with and without the contribution of the UVLF dataset.
The individual datasets provide different constraining power on the astrophysical parameters; the UVLF data is the most constraining, shrinking the posterior volume to $1.1\%$ of the prior volume, followed by CXB, CRB, and HERA at $\sim 50-52\%$ and lastly SARAS 3 at $\sim 87\%$. Note that these are individual constraints from each dataset on the full set of astrophysical parameters. Different datasets can provide more or less constraining power for specific parameters. The combination of SARAS 3, HERA, CXB and CRB show great synergies with the posterior volume corresponding to $\sim 28\%$ of the prior. With the inclusion of UVLF data, the complete joint constraints are the strongest, as expected, bringing the posterior volume further down to $0.3\%$ of the prior.
@@ 5.1 Constraints on $V_{c}$
Figure 5: Marignal 1D and 2D posteriors of SFE parameters from individual UVLF redshift bins. As expected, the lowest redshift datapoints are most constraining. Furthermore, since a single redshift bin contains no redshift-evolution information, the PDF of the $\beta_{\star}$ parameter remains flat, and the $\tilde{f}_{\star,\mathrm{II}}$ parameter compensates by shifting the peak of the 1D posterior PDF to higher values with increasing redshift.
@@
The constraints on $V_{c}$ are largely informed by the combination of SARAS 3 and UVLF data. The CXB and CRB datasets return flat 1D posterior PDFs, while HERA data has a weak (statistically insignificant) preference for lower values of $V_{c}$.
@@
The confidence levels (CLs) here, and for the rest of the results, are defined as the smallest cumulative distribution function (CDF) intervals containing the required probability volume. At $1\sigma$, this $V_{c}$ constraint alone would imply that star-formation is suppressed in most molecular-cooling halos $(V_{c}<V_{c,\text{atm}}=16.5\,\mathrm{km}\,\mathrm{s}^{-1})$. At $2\sigma$ however, the data only weakly disfavours small halos, allowing for $V_{c}\gtrsim 5.3\,\mathrm{km}\,\mathrm{s}^{-1}$, and 1D posterior PDF does not go to zero probability at the lower end. Although this has potential to be a powerful constraint, it has two added complexities: (i) we find that $V_{c}$ is degenerate with the SARAS 3 foreground parameters. In particular, the lowest order polynomial coefficients $a_{0}$ and $a_{1}$ (which have the largest effect due to the nature of the log polynomial function used) show a negative correlation with $V_{c}$. This is because (for all other astrophysical parameters fixed) lower values of $V_{c}$ correspond to deeper 21-cm absorption troughs in the SARAS 3 band. The foreground model compensates for this by assuming a larger $a_{0}$ and $a_{1}$, which boosts the foreground by tens of  mK, especially at the low redshift edge of the SARAS 3 band ($z\sim 15\,$ 2020\,2020). This is seen in the 2D posterior PDFs in Figure 13. (ii) The SARAS 3 foreground model may subsume some of the 21-cm signal, which would bias the results. This is not unlikely if the true signal lies at the low redshift end of the SARAS 3 band.
At the other end of the prior, the UV luminosity functions, by means of direct observation of faint magnitude galaxies at ${M_{\text{UV}}\lesssim-17\,\mathrm{mag}}$, provide a hard upper limit on how high $V_{c}$ can be. The data disfavour:

### v2

@@ 4.3 Joint inference
% $\begin{split}\text{\% prior volume consistent with data}&=100\times\dfrac{V_{\text{posterior}}}{V_{\text{prior}}}\\ &\approx 100\times\exp(-\mathcal{D}_{\text{KL}}).\end{split}$ (36)
@@ 5 Results and discussion
Figure 4 shows the posterior on the astrophysical parameters recovered for the different datasets independently and when they are combined in a joint analysis. A detailed comparison of constraints between the datasets, excluding UVLF observations, is done in Pochinda et al. (2024) using an SFE model that is constant in the atomic cooling regime (dashed orange line in Figure 1). Their results are retrieved in our analysis as a special case of our more general SFE model — i.e., the posterior PDF of $M_{0}$ saturates to the upper prior limit $10^{11}\,\mathrm{M}_{\odot}$, rendering $\alpha_{\star}$ and $\beta_{\star}$ parameters redundant (since most halos are $10^{11}\,\mathrm{M}_{\odot}$ at $z\geq 6$). Furthermore, the $\tilde{f}_{\star,\mathrm{II}}$ posterior PDF saturates to the lower prior limit at $10^{-4}$. This is shown in Figure 2, where we compare the joint constraints with and without (shortened to w/o) the contribution of the UVLF dataset.
The individual datasets provide different constraining power on the astrophysical parameters; the UVLF data is the most constraining, shrinking the posterior volume to $1.2\%$ of the prior volume, followed by CXB, CRB, and HERA at $\sim 50\%$ and lastly SARAS 3 at $\sim 88\%$. Note that these are individual constraints from each dataset on the full set of astrophysical parameters. Different datasets can provide more or less constraining power for specific parameters. The combination of SARAS 3, HERA, CXB and CRB show great synergies with the posterior volume corresponding to $\sim 29\%$ of the prior. With the inclusion of UVLF data, the complete joint constraints are the strongest, as expected, bringing the posterior volume further down to $0.3\%$ of the prior.
@@ 5.1 Constraints on $V_{c}$
Figure 5: Marginal 1D and 2D posteriors of SFE parameters from the full UVLF dataset and individual redshift bins. As expected, the lowest redshift bins are most constraining. Furthermore, since a single redshift bin contains no redshift-evolution information, the PDF of the $\beta_{\star}$ parameter remains flat, and the $\tilde{f}_{\star,\mathrm{II}}$ parameter compensates by shifting the peak of the 1D posterior PDF to higher values with increasing redshift in order to explain the high-$z$ JWST observations.
@@
The minimum circular virial velocity $V_{c}$ for star-formation in halos is largely informed by the combination of SARAS 3 and UVLF data. The CXB and CRB datasets return flat 1D posterior PDFs, while HERA data has a weak (statistically insignificant) preference for lower values of $V_{c}$.
The dominant constraint on $V_{c}$ comes from the UVLFs. They directly probe faint magnitude galaxies at $M_{\text{UV}}\lesssim-17\,\mathrm{mag}$}, providing a hard upper limit on how high $V_{c}$ can be. The data disfavour [^11] The confidence levels (CLs) here, and for the rest of the results, are defined as the smallest intervals in the cumulative distribution function (CDF) containing the required probability volume.:
$\log_{10}\left(V_{c}/\,\mathrm{km}\,\mathrm{s}^{-1}\right)$ $\lesssim 1.41\,(1.79)\text{ at 68\% (95\%) CL}$
$\implies V_{c}$ $\lesssim 25.7\,(62.1)\,\mathrm{km}\,\mathrm{s}^{-1}\text{ at 68\% (95\%) CL}.$
Furthermore, $V_{c}>95\,\mathrm{km}\,\mathrm{s}^{-1}$ lies outside the $5\sigma$ confidence region, which corresponds to a DM halo mass of $M_{h}\approx 3\times 10^{10}\,\mathrm{M}_{\odot}$ at $z=6$, and stellar masses of $M_{\star}\sim 10^{7}\,\mathrm{M}_{\odot}$ in galaxies (assuming percent-level star-formation efficiency) which is consistent with the faintest galaxies observed at $z\sim 6$ (see, e.g., Figure 5 of Simmonds et al., 2024). This constraint is robust across different redshift data, as seen in Figure 5 (first column). In order to further tighten this derived constraint, future works will include lensed luminosity functions such as the ones in Bouwens et al. (2022) that go down to $M_{\star}\sim 10^{5-6}\,\mathrm{M}_{\odot}$ at $z\sim 6$ (Bouwens et al., 2021b). Since our model has the added benefit of including photoheating feedback at the faint end, these lensed UVLFs can shed light on the suppression of star-formation in small halos.
Recently, Sipple & Lidz (2024) performed an analysis of lensed UVLF data using truncated SFE models and found that non-lensed UVLFs disfavour a truncation of star-formation above $M_{h}\sim 3\times 10^{10}\,\mathrm{M}_{\odot}$ at $z=6$ (see their Figure 11), consistent with our results. Furthermore, they suggest that lensed UVLFs disfavour a truncation above $2\times 10^{9}\,\mathrm{M}_{\odot}$ at $z=6$ (i.e., strong photoheating feedback disfavoured) at $2\sigma$ confidence. This would correspond to a $V_{c}\lesssim 36.5\,\mathrm{km}\,\mathrm{s}^{-1}$ being favoured at $2\sigma$, indeed stronger than our $V_{c}\lesssim 62.1\,\mathrm{km}\,\mathrm{s}^{-1}$ constraint. Another recent multi-wavelength analysis by Zvi Katz et al. (2025) similar to ours also finds that UVLFs play a crucial role in providing an upper bound on the minimum halo mass required for star-formation.
[…]
$\log_{10}\left(V_{c}/\,\mathrm{km}\,\mathrm{s}^{-1}\right)$ $\gtrsim 1.16\,(0.72)\text{ at 68\% (95\%) CL}$
$\implies V_{c}$ $\gtrsim 14.4\,(5.3)\,\mathrm{km}\,\mathrm{s}^{-1}\text{ at 68\% (95\%) CL}.$
At $1\sigma$, this $V_{c}$ constraint alone would imply that star-formation is suppressed in most molecular-cooling halos $(V_{c}<16.5\,\mathrm{km}\,\mathrm{s}^{-1})$. At $2\sigma$ however, the data only weakly disfavours small halos, allowing for $V_{c}\gtrsim 5.3\,\mathrm{km}\,\mathrm{s}^{-1}$, and 1D posterior PDF does not go to zero probability at the lower end. Although this has potential to be a powerful constraint, it has two added complexities: (i) we find that $V_{c}$ is degenerate with the SARAS 3 foreground parameters. In particular, the lowest order polynomial coefficients $a_{0}$ and $a_{1}$ (which have the largest effect due to the nature of the log polynomial function used) show a negative correlation with $V_{c}$. This is because (for all other astrophysical parameters fixed) lower values of $V_{c}$ correspond to deeper 21-cm absorption troughs in the SARAS 3 band. The foreground model compensates for this by assuming a larger $a_{0}$ and $a_{1}$, which boosts the foreground by tens of  mK, especially at the low redshift edge of the SARAS 3 band ($z\sim 15$). This is seen in the 2D posterior PDFs in Figure 4. (ii) The SARAS 3 foreground model likely subsumes some of the 21-cm signal. Thus, in future works, it would be good to check if the choice of the form of foreground model biases our results.
In the joint analysis combining all datasets, the inclusion of SARAS 3 data weakens the upper limit on $V_{c}$ set by UVLFs as shown in Figure 3. The joint constraints yield a $V_{c}$ value of:
[…]
$V_{c}$ $\lesssim 25.2\,(62.0)\,\mathrm{km}\,\mathrm{s}^{-1}\text{ at 68\% (95\%) CL, with SARAS~3.}$
Table 3 summarizes these constraints at both 68% and 95% confidence levels, converted to $T_{\text{vir}}$ and $M_{\text{vir}}(z=6)$ for ease of comparison with works that use alternate parameterizations. Given the SARAS 3 constraint’s sensitivity to the foreground fit, even a marginal change in the foreground model can lead to stronger exclusion of low $V_{c}$ values. This is seen in Figure 6 (first panel) in the case of joint constraints without high-$z$ UVLF data. In the absence of a more physically motivated foreground model, we refrain from over-interpreting this result but emphasize the potential for future 21-cm signal measurements in providing constraints on $V_{c}$.
Our results demonstrate the synergy that exists between the 21-cm signal data and high-$z$ galaxy surveys from HST/JWST, even when the former is a non-detection, in constraining the star-forming properties of small halos.
5.2 Constraints on the star-formation efficiency
The 21-cm signal data is not informative enough yet to constrain the SFE parameters independently and, in the absence of the UVLF data, the 1D posterior PDF for the scaling of the SFE $\tilde{f}_{\star,\mathrm{II}}$ saturates to the lower limit of the prior at $10^{-4}$, while the probability of the turning point for the SFE power law slope $M_{0}$ saturates to the upper limit $\sim 10^{11}\,\mathrm{M}_{\odot}$ (see Figure 2). This is consistent with predictions from a fixed constant SFE model as adopted in Pochinda et al. (2024). Thus, constraints on the star-formation efficiency parameters $\{\tilde{f}_{\star,\mathrm{II}},M_{0},\alpha_{\star},\beta_{\star}\}$ are almost entirely driven by UVLF data. Table 4 summarizes our results at 68% and 95% confidence levels. As seen in the 2D posterior plot for $\tilde{f}_{\star,\mathrm{II}}$ and $M_{0}$ in Figure 4, there is a strong degeneracy below ${M_{0}}\lesssim 10^{10.5}\,\mathrm{M}_{\odot}$ which is well-described by a simple linear fit to the 2D histogram:
$\log_{10}\left(M_{0}/\,\mathrm{M}_{\odot}\right)\approx 0.9\times\log_{10}\left(\tilde{f}_{\star,\mathrm{II}}\right)+12,$ (37)
before saturating at $\tilde{f}_{\star,\mathrm{II}}<0.1$. This is as expected from the observational limit on faint galaxies, as discussed in the previous section. To better understand the constraining power of the UVLF data, we now perform different redshift cuts and analyze the SFE parameters inferred.
Figure 6: Marginal 1D posteriors of the SFE parameters $\tilde{f}_{\star,\mathrm{II}}$, $M_{0}$, $\alpha_{\star}$, and $\beta_{\star}$ from all observational datasets, with the UVLFs split into two redshift bins: z=7z=7\, 10\,10 as the ‘low-$z$’ (purple) and z=11z=11\, 14.5\,14.5 as the ‘high-$z$’ (pink) fits, with the $z=6$ data points from B21 included in both (see Footnote 12). The main takeaway here is that the redshift-evolution term $\beta_{\star}$ showing little to no evolution at low-$z$ (i.e., consistent with 0) and a rapid evolution at high-$z$. The $V_{c}$ constraints, discussed in Section 5.1, show slight variation depending on the SARAS 3 foreground parameter, with the low-$z$ fit exhibiting the strongest synergy between the UVLF and SARAS 3 data.
Parameter Prior Joint w/o high-$z$ UVLF Joint w/o low-$z$ UVLF Joint
(low-$z$ fit) (high-$z$ fit) (joint fit)
@@ 5.2 Constraints on the star-formation efficiency
[…]
5.2.1 How much do individual redshifts contribute?
We first quantify the constraining power of the UVLF data from different redshift bins individually. This is shown in Figure 5. The SFE parameter space is constrained to $\lesssim 10\%$ of the prior volume by data at $z\approx 6\,$ 9\,9, $\lesssim 20\%$ at $z\approx 10\,$ 13\,13, and $\sim 60\%$ at $z\approx 14.5$ as the data becomes increasingly uncertain at high redshifts.
The constraint on $V_{c}$ is consistent across the redshift bins. In particular, observations of faint galaxies at $z=12.5$ provide just as good constraints as those at $z=6$. Deeper observations in fainter magnitude bins, including lensed objects with JWST, are expected to provide more stringent upper limits on $V_{c}$ in the near future.
Similarly, the constraints on the halo-mass dependence of the SFE, $\alpha_{\star}$, also show consistency between redshift bins. The scatter around the mean value inferred from the combined datasets is well within errors. Furthermore, since there is no redshift-evolution information encoded in individual redshift bins, the PDF of $\beta_{\star}$ is flat across all redshifts. To compensate for this lack of evolution, the peak of the PDF for $\tilde{f}_{\star,\mathrm{II}}$ shifts to higher values with increasing redshift, thus hinting at an enhanced SFE at early times.
5.2.2 How does the SFE evolve with redshift?
In order to properly quantify the evolution of the SFE with redshift, $f_{\star,\mathrm{II}}\propto\left[(1+z)/7\right]^{-\beta_{\star}}$, we now take a closer look at the $\beta_{\star}$ parameter. We split the UVLF dataset into two redshift bins, now including the rest of the observational datasets (SARAS 3, HERA, CXB and CRB data) for both, to compare to the joint constraint which uses the complete UVLF dataset across all redshifts. We denote z=7z=7\, 10\,10 as the ‘low-$z$’ cut and z=11z=11\, 14.5\,14.5 as the ‘high-$z$’ cut, including the $z=6$ data point from B21 in both. [^12] This ‘anchoring’ at $z=6$ is done for two reasons. First, the SFE as defined in Equation 7 and 8, is normalized to this redshift which means the $\beta_{\star}$ parameter has no effect at $z=6$. Secondly, by fixing $\tilde{f}_{\star,\mathrm{II}}$, $\alpha_{\star}$ and $M_{0}$ to the best-fit values at $z=6$, we can isolate the effect of SFE evolution via $\beta_{\star}$. The choice of redshift bins defines the high-redshift frontier before and after the launch of the JWST, hence informing us on how the new instrument has changed our understanding of the SFE. Figure 6 shows the result of this analysis. The $\tilde{f}_{\star,\mathrm{II}}$ and $\alpha_{\star}$ parameters are consistent across the two redshift ranges and $M_{0}$ remains degenerate with $\tilde{f}_{\star,\mathrm{II}}$, as suggested by Equation 37. However, $\beta_{\star}$ shows a stark contrast:
  • Joint w/o high-$z$ UVLF (or low-$z$ fit for short): At $z\leq 10$, the UVLF data favours a very slow/non-evolving SFE. The 1D posterior PDF of $\beta_{\star}$ saturates at the lower edge of the prior, favouring $\lesssim 1.4$ at 68% CL, and remains consistent with $\sim 0$. This behaviour aligns with pre-JWST inferences using HST data (Tacchella et al., 2013; Mason et al., 2015; Bouwens et al., 2015; Tacchella et al., 2018; Harikane et al., 2018; Bouwens et al., 2021a; Stefanon et al., 2021), where a simple redshift-indepenent parameterization calibrated at $z\lesssim 4$ can reproduce the observed UVLFs at higher redshifts. This trend of slow evolution is also consistent up to $z\approx 10$ with the FirstLight numerical simulation suite (Ceverino et al., 2024).
  • Joint w/o low-$z$ UVLF (or high-$z$ fit for short): At $z>10$, the UVLF data favours a rapidly evolving SFE with the 1D posterior PDF of $\beta_{\star}$ having a weighted mean of $\langle\beta_{\star}\rangle=3.6$. This deviation from the low-$z$ behaviour is an interesting outcome of the abundance of bright galaxies observed in the JWST data. The bright excess has been alternatively attributed in the literature to either a top-heavy IMF of the early populations of stars (Inayoshi et al., 2022; Wang et al., 2023; Trinca et al., 2024; Hutter et al., 2025), or stochasticity/variability in star-forming galaxies (Mason et al., 2023; Mirocha & Furlanetto, 2023; Sun et al., 2023a, b; Pallottini & Ferrara, 2023; Shen et al., 2023). [^13] For completeness, we note that other solutions have also been proposed. This includes the presence of a hidden population of AGN (Hegde et al., 2024), or simply that a formation epoch of $z\approx 15$ implies young, bright stellar population (requiring no new physics to explain current observations; Donnan et al., 2025). However, both these solutions give rise to other problems — for example, Cueto et al. (2024) find that a top-heavy IMF alone leads to strong stellar feedback which inhibits stellar mass growth. Similarly, Nikolić et al. (2024) and Yung et al. (2024) conclude that larger than physical scatter would be needed to explain the bright galaxies. Although a redshift dependence in the SFE has been previously proposed (e.g. Sun & Furlanetto, 2016), we constrain this SFE evolution for the first time at high-$z$ using a sub-grid analytic prescription. The redshift evolution is quantitatively similar to the feedback free burst (FFB) galaxies explored in Dekel et al. (2023); Li et al. (2024), as well as the density-modulated SFEs (DMSFE) in Somerville et al. (2025). Indeed, we shall see how our inferred SFE posteriors align with the former in the next section. A similar transition from low SFE to high SFE at $z\sim 10$ has also been inferred in Chakraborty & Choudhury (2025, see their Figure 10).
  • Joint (or joint fit): The joint constraints, combining both low-$z$ and high-$z$ UVLF data, give a weighted mean of $\langle\beta_{\star}\rangle=2.9$. This fit is effectively a compromise between the two redshift bins, showing a moderate evolution of the SFE. Although not an ideal fit at the lower redshifts due to small error-bars on the data, it provides a simple parameterization that can be applied across the entire redshift range.
Figure 7: Functional posteriors of the UVLF (left) and constrained SFE (right) for the fit using UVLFs at low and high redshift (including all other datasets). In the upper three panels, $z=6,8,10$ we use the low-$z$ fit, while in the lower two panels, $z=12.5,14.5$ we use the high-$z$ fit (see Figure 6 and Table 4). In all the panels, the filled coloured contours show $1,2,3\sigma$ confidence of the functional posterior (also showing functional SFE priors in the right panels in grey contours), while the thick black line shows the posterior-mean model (i.e., the weighted means listed in Table 4). Note that the black line and contours need not align since $\langle f(x|\theta_{\text{ast}})\rangle\neq f(x|\langle\theta_{\text{ast}}\rangle)$. The UVLF data (scatter points, left panels) used for the constraints are shown as a cross-check to see whether the model fits the data. In the right column, we compare the SFE posteriors to various models in the literature: halo-abundance matching results from Sun & Furlanetto (2016, SF16), empirical fits from Tacchella et al. (2018, T+18), UniverseMachine (Behroozi et al., 2019, B+19), semi-empirical model of Sipple & Lidz (2024, SL24), feedback-free burst model of Dekel et al. (2023, L+24); Li et al. (2024, L+24), and HST/Spitzer based results from Stefanon et al. (2021, S+21).
Figure 8: Evolution of SFE at fixed halo masses $M_{h}=10^{9}\,\mathrm{M}_{\odot}$ (left, triangles) and $M_{h}=10^{10}\,\mathrm{M}_{\odot}$ (right, circles) with redshift, for the case of the three joint analyses: low-$z$ UVLFs (purple) and high-$z$ UVLFs (pink), separated by a cut at $z\approx 10.5$ (dashed vertical line), and all UVLFs (black). Unfilled markers and dashed coloured lines denote extrapolations outside of the redshift range of the fits. In the case of $M_{h}=10^{10}\,\mathrm{M}_{\odot}$, the error bars are $1\sigma$ confidence levels while for $M_{h}=10^{9}\,\mathrm{M}_{\odot}$, the SFE are $1\sigma$ upper limits and relatively unconstrained due to the lack of UVLF data at the faint magnitude end. See Figure 7 for the full SFE posteriors.
Figure 9: Evolution of the SFRD, calculated by integrating the UVLF ($\dot{\rho}_{\star,\mathrm{II}}=\int\Phi\dot{M}_{\star,\mathrm{II}}dM_{\text{UV}}$}) in the magnitude range $M_{\text{UV}}\in[-23,-17]$. The filled contours show functional posteriors of the joint fit, while the lines show the posterior-mean model (see Table 4) for the joint fit (black), low-$z$ fit (purple) and high-$z$ fit (pink). The lines are solid in the redshift range of the fit and dashed outside where the model is extrapolated. Various literature values are shown for comparison: results from HST in blue (Bouwens et al., 2015; Finkelstein et al., 2015; McLeod et al., 2016; Bhatawdekar et al., 2019) and JWST in green (Harikane et al., 2023a; Donnan et al., 2024; Finkelstein et al., 2024; Whitler et al., 2025; Pérez-González et al., 2025).
5.2.3 SFE functional posteriors and best fit
As suggested in the previous section, the star-formation efficiency appears to have a complex evolution with time. Namely, the $\beta_{\star}$ parameter alone is not enough to capture the evolution at both low-$z$ and high-$z$. Thus, our inference across the full redshift range z=6z=6\, 14.5\,14.5 is essentially a composite of the low-$z$ fit at $z\leq 10$ and the high-$z$ fit at $z>10$. In order to visualize the results from our Bayesian analysis in the two redshift bins, we use fgivenx (Handley, 2018) to convert our parameter posterior samples into functional posteriors of the SFE and its associated UVLF (as a cross-check to see how our model fits the data). This is shown in Figure 7.
We also plot the posterior-mean model in the two redshift bins (Table 4 using Equation 7) for comparison, alongside other empirical SFE models from literature. Note that other works may use different cosmologies and IMFs, with potentially different halo mass definitions and mass accretion models which could lead to biases compared to our results. Hence, renormalizing their results to our choice of cosmology and SFR-luminosity conversion is non-trivial and we show the SFE curves as presented in their works.
At a glance, one can see that the $1\sigma$ SFE constraints are generally informative up to $M_{h}\gtrsim 10^{10}\,\mathrm{M}_{\odot}$ at $z=6$ moving down to $M_{h}\gtrsim 10^{9}\,\mathrm{M}_{\odot}$ at $z=14.5$. At lower halo masses, the SFE is unconstrained and hence, the degeneracy between $\tilde{f}_{\star,\mathrm{II}}$ and $M_{0}$ allows for a wide range of curves that flatten below the observed mass threshold. The shrinking of SFE priors (grey contours in Figure 7) to posteriors (coloured contours in the same Figure) at the low mass end is due to our implicit assertion that the SFE is continuous across halo masses, which is not a priori guaranteed, but a reasonable assumption (but see also the ‘bursty’ or ‘high scatter’ model of Sipple & Lidz, 2024, in their Figure 11). Furthermore, the steep contours for SFE and the drop in UVLF at the faint magnitude end (which is redshift dependent) are a physical feature — this is the direct result of the constraint on high values of minimum circular velocity $V_{c}$ (equivalently $T_{\text{vir}}$ or $M_{\text{vir}}$) for star-formation in DM halos.
At the lowest redshift, $z=6$, we find strong agreement with the SFE magnitude and slope of Tacchella et al. (2018, T+18) and Sun & Furlanetto (2016, SF16) to $\lesssim 2\sigma$, both of which predict $f_{\star,\mathrm{II}}\propto M_{h}$ (or equivalently, $M_{\star}\propto M_{h}^{2}$). The former is an empirical redshift-independent model based on informed accretion histories from $N$-body simulations, while the latter uses halo-abundance matching assuming smooth gas accretion. The results are also within $2\sigma$ of the deepest Spitzer and HST-calibrated results from Stefanon et al. (2021, S+21). Compared to the recent work of Sipple & Lidz (2024, SL24, in particular their ‘B21+B22’ model), although roughly consistent at $3\sigma$, we find a steeper slope than their $f_{\star,\mathrm{II}}\propto M_{h}^{0.56}$ and lower efficiency at the the faint end. The feedback-free burst model from Dekel et al. (2023, L+24, their $\epsilon_{\text{max}}=0.2$ model); Li et al. (2024, L+24, their $\epsilon_{\text{max}}=0.2$ model) is a factor of few lower than our prediction, which is expected because FFB galaxies are less abundant compared to non-FFB galaxies at this redshift.
At $z=8$, we find an almost exact match with the SFE of T+18 and SF16. We are also $\lesssim 2\sigma$ consistent with L+24 across the halo mass range, and with SL24 at the high mass end (deviating by a factor of few at the low mass end). The UVLF best-fit model at this redshift is in fact more consistent with latest spectroscopic measurements from JWST (Rojas-Ruiz et al., 2025), lying slightly above the B21 data points at the bright end.
At $z=10$, we are consistent with most works at the low mass end, while predicting an intermediate SFE at the high mass end, between the FFB model of L+24 and others. At $z>10$, we predict an order of magnitude higher SFEs than T+18 and strongly favour the FFB model of L+24 quantitatively.
This result recovers the slow evolution of the SFE inferred in pre-JWST HST-calibrated results at $z\lesssim 10$, and enhanced SFE from a rapid evolution at $z\gtrsim 10$ in post-JWST models (a transition consistent with the findings of Chakraborty & Choudhury 2025, although with a different parameterization). Across the entire redshift range, the SFE slope $\langle\alpha_{\star}\rangle\approx 1.1$ is a reasonable fit to data (see Table 4 for exact values and error-bars), which is consistent with T+18 within errors.